OpenStudy (anonymous):

Find the distance between (-2,-1) (7,y) is 15

OpenStudy (mertsj):

That doesn't make sense.

OpenStudy (anonymous):

find y if the distance between the pionts a(-2,-1) b(7,y) is 15?

OpenStudy (mertsj):

Use the distance formula.

OpenStudy (mertsj):

\[\sqrt{(7-(-2))^{2}+(y-(-1))^{2}} = 15\]

OpenStudy (mertsj):

\[\sqrt{81+y ^{2}+2y+1}^{ } = 15\]

OpenStudy (mertsj):

\[y ^{2}+2y+82 = 225\]

OpenStudy (mertsj):

\[y ^{2}+2y -143 = 0\]

OpenStudy (mertsj):

\[(y+13)(y-11)= 0\]

OpenStudy (mertsj):

y = -13 or y = 11